# OmniArch: Building the Foundation Model for Scientific Computing

### v2

@@ 5 Conclusion
Appendix
@@ Appendix A Table of notations
$\sigma_{u}$ Variance of physical field $u$
$L^{u}_{\text{sim}}$ nRMSE loss of physical field $u$
$L_{\text{sim}}$ nRMSE loss of all physical fields
@@ Appendix D Baseline implementation details
In our experiments, we adopt the benchmarking framework provided by PDEBench  [34] and select three well-established methods for comparative analysis. Furthermore, we have incorporated the Multiple Physics Pre-training (MPP) model into our comparative analysis to address the need for retraining that is inherent to the aforementioned methods when faced with novel sets of conditions, the detailed training hyperparameters of FNO, U-Net, and PINN is provided in Table 6, following PDEbench [34]. The first hyperparameter of U-Net is the unroll steps (denoted as us), and the second is the train steps (denoted as ts). The hyperparameters shared by both FNO and U-Net are the initial steps (denoted as is) and batch size (denoted as bs). The hyperparameter in PINNs is the hidden size (denoted as hid). The learning rate, shared by FNO, U-Net, and PINNs, is denoted as lr.
@@
Physics-Informed Neural Networks (PINNs)  [26]. PINNs utilize neural networks to solve differential equations by embedding physical laws into a multi-objective optimization framework, minimizing PDE residuals and boundary/initial condition errors  [5].
U-Net  [29]. U-Net, designed for biomedical image segmentation, uses an encoder-decoder structure for context capture and precise localization  [31, 7]. We adapt U-Net into 1D and 3D forms to analyze spatio-temporal patterns in physical fields.
Fourier Neural Operator (FNO)  [12]. FNO pioneers in learning function-to-solution mappings by parameterizing integral kernels in the Fourier domain, enabling efficient and accurate resolution-invariant neural operators.
PDEformer-1  [39]. PDEformer-1 is a neural solver capable of simultaneously addressing various types of 1D partial differential equations. It uses a graph Transformer and implicit neural representation (INR) to generate mesh-free predicted solutions.
Multiple Physics Pre-training (MPP)  [17]. MPP extends PDEBench’s 2D physics scenarios to learn versatile features for predicting dynamics across various physical systems and comprises pre-training and fine-tuning phases, warranting its inclusion in our comparative analysis.
ORCA-SWIN  [30, 13]. ORCA fine-tunes the SWIN Transformer for different PDEs by first aligning the embedded feature distribution of the target PDE data with the pre-training modality, and then refining the model on this aligned data to effectively leverage shared knowledge across various PDEs.
@@ F.3 Pre-training process of PDE-Aligner
In our architecture, the PDE-Aligner is divided into two components: a text encoder and a physics encoder. The text encoder utilizes the pre-trained albert-math model [28], which is highly capable of processing LaTeX-encoded PDE captions due to its extensive training on a large corpus of LaTeX data. For the physics encoder, we employ the pre-trained Fourier encoder from OmniArch, known for its strong ability to capture physical field features. We adopt a large-batch contrastive learning approach similar to SimCLR [4]. The training involves a stochastic sampling strategy with an equal probability (50%) of selecting either canonical PDE captions sourced directly from textbooks or augmented PDE captions. The latter is assumed to enhance the text encoder’s generalization capabilities while retaining critical PDE information in textual form. The weights of the text encoder and physics encoder are fixed during the PDE-Aligner training process. The training data details for PDE-Aligner are shown in Table 9, and the hyperparameter settings are provided in Table 10.
@@ H.3 Fine-tuned for Inverse Problems
Demonstrating a model’s capability to infer hidden physical parameters from known equations is a critical test of its ability to learn underlying physics. Following the methodology of MPP [17], we evaluate our model on two inverse problems for incompressible Navier-Stokes equations: 1) Forcing Identification, and 2) Buoyancy Determination.
@@ H.7 GPU Memory Usage and Inference Time
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The motivation and contribution are shown in the introduction, and the results match the claim.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations is discussed in section B.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when the image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to addressing problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: In the design of OmniArch(see section 3.1), we provide theoritical analysis.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in the appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We include very detailed experiment setup C, data cook receipts and the training detailsE in appendix. Also, we provide the code in the supplementary files.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide full code and running scripts in the supplementary. The data we used include PDEBench(https://github.com/pdebench/PDEBench) and PDEAreana(https://pdearena.github.io/pdearena/) are all open-sourced.
  25. Guidelines:
    • The answer NA means that the paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide the details of data preparation, baselines, software/hardware settings, metrics, and other details in section 4 and appendix C, appendix E.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in the appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We use the average performance of repeated experiments in the main results, and show the standard deviation.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar and then state that they have a 96% CI if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We include our hardware resources and training details in the Appendix E.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We use the correct and broad domain conflicts to preserve anonymity.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: There is no societal impact of the work performed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for the responsible release of data or models that have a high risk for misuse (e.g., pre-trained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not have such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make the best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited, and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The dataset and models we used in 4 is properly cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We have included the correct license in the supplementary.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v3

@@ 5 Conclusion
Acknowledgement
This work was supported by the National Science and Technology Major Project(No.2022ZD0117800), and Young Elite Scientists Sponsorship Program by CAST(No.2023QNRC001). This work was also sponsored by CAAI-Huawei MindSpore Open Fund (CAAIXSJLJJ2023MindSpore12) and developed on openl community. Thanks for the computing infrastructure provided by Beijing Advanced Innovation Center for Big Data and Brain Computing.
Impact Statement
OmniArch represents a significant advancement in scientific computing. It unifies multi-scale and multi-physics PDE-solving capabilities within a single foundation model framework. This unified approach has profound implications for accelerating scientific discovery and engineering applications across domains such as fluid dynamics, weather forecasting, and materials science. The model’s demonstrated ability to handle diverse physical systems and grid resolutions while maintaining physical consistency could dramatically reduce the computational resources required for complex simulations in industrial and research settings. Additionally, there are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Supplementary Material for:
OmniArch: Building Foundation Model For Scientific Computing
Contents
  • A.
    Table of notations A
  • B.
    Limitations B
  • C.
    Dataset details C
    • C.1.
      OmniArch Pre-training Dataset C.1
    • C.2.
      Dataset For Zero-shot Learning C.2
  • D.
    Baseline implementation details D
  • E.
    OmniArch implementation details E
    • E.1.
      Pre-training OmniArch E.1
    • E.2.
      Fine-tuning OmniArch E.2
    • E.3.
      Parameter Efficiency Analysis E.3
  • F.
    PDE-Aligner implementation details F
    • F.1.
      PDE-Aligner Pre-training Dataset F.1
    • F.2.
      Examples of Generated PDEs F.2
    • F.3.
      Pre-training process of PDE-Aligner F.3
  • G.
    Further ablation study G
    • G.1.
      Dynamic Prompt Length for Efficient Inference H.2
    • G.2.
      Fine-tuned for Inverse Problems H.3
  • H.
    More results H
    • H.1.
      Zero-shot Learning Capability H.1
    • H.2.
      Rollout Predictions H.4
    • H.3.
      Multi-scale Inference Results H.5
    • H.4.
      More results in different problem settings H.6
    • H.5.
      GPU Memory Usage and Inference Time H.7
    • H.6.
      Comparison with Traditional Solvers H.8
  • I.
    More Discussions I
    • I.1.
      Meta-Learning vs. Scaling Laws in PDE SolvingI.1
@@ Appendix A Table of notations
$\sigma_{u}$ Normalization factor $\|u\|_{2}^{2}+\epsilon$ for nRMSE calculation
$L^{u}_{\text{sim}}$ Normalized RMSE loss: $\sqrt{\mathbb{E}[(u^{\text{pred}}-u)^{2}]/\sigma_{u}}$
$L_{\text{sim}}$ Mean nRMSE across all physical fields
@@ Appendix D Baseline implementation details
In our experiments, we adopt the benchmarking framework provided by PDEBench  (Takamoto et al., 2022) and select three well-established methods for comparative analysis. Furthermore, we have incorporated the Multiple Physics Pre-training (MPP) model into our comparative analysis to address the need for retraining that is inherent to the aforementioned methods when faced with novel sets of conditions, the detailed training hyperparameters of FNO, U-Net, and PINN is provided in Table 8, following PDEbench (Takamoto et al., 2022). The first hyperparameter of U-Net is the unroll steps (denoted as us), and the second is the train steps (denoted as ts). The hyperparameters shared by both FNO and U-Net are the initial steps (denoted as is) and batch size (denoted as bs). The hyperparameter in PINNs is the hidden size (denoted as hid). The learning rate, shared by FNO, U-Net, and PINNs, is denoted as lr.
@@
Physics-Informed Neural Networks (PINNs)  (Raissi et al., 2019). PINNs utilize neural networks to solve differential equations by embedding physical laws into a multi-objective optimization framework, minimizing PDE residuals and boundary/initial condition errors  (Cuomo et al., 2022).
U-Net  (Ronneberger et al., 2015). U-Net, designed for biomedical image segmentation, uses an encoder-decoder structure for context capture and precise localization  (Siddique et al., 2021; Du et al., 2020). We adapt U-Net into 1D and 3D forms to analyze spatio-temporal patterns in physical fields.
Fourier Neural Operator (FNO)  (Li et al., 2020). FNO pioneers in learning function-to-solution mappings by parameterizing integral kernels in the Fourier domain, enabling efficient and accurate resolution-invariant neural operators.
PDEformer-1  (Ye et al., 2024). PDEformer-1 is a neural solver capable of simultaneously addressing various types of 1D partial differential equations. It uses a graph Transformer and implicit neural representation (INR) to generate mesh-free predicted solutions.
Multiple Physics Pre-training (MPP)  (McCabe et al., 2023). MPP extends PDEBench’s 2D physics scenarios to learn versatile features for predicting dynamics across various physical systems and comprises pre-training and fine-tuning phases, warranting its inclusion in our comparative analysis.
ORCA-SWIN  (Shen et al., 2023; Liu et al., 2021). ORCA fine-tunes the SWIN Transformer for different PDEs by first aligning the embedded feature distribution of the target PDE data with the pre-training modality, and then refining the model on this aligned data to effectively leverage shared knowledge across various PDEs.
@@ E.2 Fine-tuning OmniArch
E.3 Parameter Efficiency Analysis
Table 11: Static Parameter Distribution (Millions)
Model Component OmniArch-B (316M) OmniArch-L (672M)
Shared Backbone 138 (43.7%) 435 (64.7%)
1D Encoder/Decoder 0.3 0.4
2D Encoder/Decoder 7.0 9.0
3D Encoder/Decoder 171 227
Table 12: Active Parameters During Task Execution (Millions)
Model PDE Type
1D PDEs 2D PDEs 3D PDEs
OmniArch-B 138 144 308
OmniArch-L 435 445 663
As illustrated in Table 11 and Table 12, the parameter distribution reveals OmniArch’s hierarchical design philosophy. Three key observations emerge: (1) The shared backbone dominates the parameter count (43.7–64.7%), facilitating cross-dimensional knowledge transfer while requiring only modest modality-specific additions (0.3–227M). (2) For 2D tasks (MPP’s primary domain), OmniArch-B activates merely 144M parameters—a 24.1% increase over MPP-B’s 116M that brings three key advantages: (a) unified architecture reduces system complexity, (b) enables latent cross-modal learning, and (c) provides future-proof extensibility. (3) The scaling pattern shows intelligent allocation—3D processing requires 2.1–2.7× more dedicated parameters than 2D, reflecting its inherent higher dimensionality while maintaining efficient reuse of the shared backbone.
@@ F.3 Pre-training process of PDE-Aligner
In our architecture, the PDE-Aligner is divided into two components: a text encoder and a physics encoder. The text encoder utilizes the pre-trained albert-math model (Reusch et al., 2022), which is highly capable of processing LaTeX-encoded PDE captions due to its extensive training on a large corpus of LaTeX data. For the physics encoder, we employ the pre-trained Fourier encoder from OmniArch, known for its strong ability to capture physical field features. We adopt a large-batch contrastive learning approach similar to SimCLR (Chen et al., 2020). The training involves a stochastic sampling strategy with an equal probability (50%) of selecting either canonical PDE captions sourced directly from textbooks or augmented PDE captions. The latter is assumed to enhance the text encoder’s generalization capabilities while retaining critical PDE information in textual form. The weights of the text encoder and physics encoder are fixed during the PDE-Aligner training process. The training data details for PDE-Aligner are shown in Table 13, and the hyperparameter settings are provided in Table 14.
@@ H.3 Fine-tuned for Inverse Problems
Demonstrating a model’s capability to infer hidden physical parameters from known equations is a critical test of its ability to learn underlying physics. Following the methodology of MPP (McCabe et al., 2023), we evaluate our model on two inverse problems for incompressible Navier-Stokes equations: 1) Forcing Identification, and 2) Buoyancy Determination.
@@ H.7 GPU Memory Usage and Inference Time
H.8 Comparison with Traditional Solvers
Table 19: Computational Efficiency Comparison (2D Advection)
Resolution FDM Time/Step (ms) OmniArch Time (ms) Speedup Relative Error
64$\times$64 1.123 23.567 0.048$\times$ 1.24$\times$
128$\times$128 15.264 23.820 0.641$\times$ 1.18$\times$
192$\times$192 75.360 24.098 3.128$\times$ 1.15$\times$
256$\times$256 254.027 24.083 10.55$\times$ 1.12$\times$
320$\times$320 583.218 23.866 24.44$\times$ 1.09$\times$
384$\times$384 1130.561 23.453 48.20$\times$ 1.07$\times$
448$\times$448 2272.206 23.677 95.96$\times$ 1.05$\times$
512$\times$512 4073.472 26.212 155.4$\times$ 1.03$\times$
Our benchmarks reveal three key advantages of OmniArch over traditional solvers:
  • Resolution Invariance: While FDM computation time scales quadratically ($O(n^{2})$) with grid resolution, OmniArch maintains nearly constant inference time (23-26ms) due to its fixed-frequency processing in the spectral domain. This yields exponential speedup (155$\times$ at 512$\times$512) for high-resolution simulations.
  • Accuracy Preservation: Despite dramatic speed improvements, OmniArch maintains comparable accuracy with relative error consistently below 1.25$\times$ of FDM results. The error margin decreases at higher resolutions (1.03$\times$ at 512$\times$512), suggesting better performance in practical high-fidelity scenarios.
  • Generalization Capability: Unlike traditional methods requiring re-discretization for new PDEs, OmniArch’s unified architecture achieves this performance across multiple physics domains (Navier-Stokes, Advection-Diffusion, etc.) without algorithmic modifications, as demonstrated in Section 4.2.
The results validate our design choice of spectral-domain processing - while sacrificing some interpretability inherent to mesh-based methods, OmniArch gains orders-of-magnitude efficiency improvements crucial for large-scale multi-physics simulations. This trade-off aligns with emerging trends in scientific ML where learned simulators complement (rather than replace) traditional methods for specific high-throughput applications.
Appendix I More Discussions
I.1 Meta-Learning vs. Scaling Laws in PDE Solving
While meta-learning methods (Chen et al., 2022; Huang et al., 2022; Cho et al., 2023) address generalization through gradient-based adaptation, OmniArch explores an orthogonal axis: scaling laws for in-context learning. The distinction mirrors “learning to optimize" versus “learning from data" paradigms—meta-PINNs refine their optimization trajectory for new PDEs, whereas foundation models leverage scale to discover physics-aware primitives. These approaches need not compete; future work might hybridize them. We may imagine meta-learning the hypernetworks of a foundation model.